L^{A}TEX News

Issue 15, December 2003

Anniversary release

Yes, it’s now 10 years since the first release in this series and, for Knuthists, this release also contains Issue 16 !

Meanwhile this Issue 15 describes the major new features in the current release whilst Issue 16 looks a little way into the future of L^{A}TEX.

LPPL – new version

Most importantly, there is now a new version, 1.3, of the L^{A}TEX Project Public Licence. Many of you will be thrilled to know that, following the exchange of over 1600 e-mail messages dissecting various aspects of its philosophy such as ‘how many angels can appear in the name of a file before it becomes non-free’, this version is now officially a DFSG (Debian Free Software Guidelines) approved license. The discussions start at

http://lists.debian.org/debian-legal/2002/
debian-legal-200207/threads.html with high

traffic throughout August to October 2002 and further heated discussions starting in April 2003 and concluding around June at

http://lists.debian.org/debian-legal/2003/
debian-legal-200306/msg00206.html.

The important features of the new version are useful clarifications in the wording, and revised procedures for making a change to the Current Maintainer of a package. Special thanks to all those people from Debian Legal who worked constructively with us on this onerous task, especially but not exclusively Jeff Licquia and Branden Robinson.

Small updates to varioref

The English has been corrected in \reftextbefore (an incompatible change). There are other extensions such

as \labelformat, \Ref, \Vref and \vpagerefnum.

Some Dutch text has also been changed and two new options added: slovak and slovene.

New and more robust commands

Many of the math mode commands for compound symbols have been made robust and a new robust command has been added: \nobreakdashes. This last is a low-level command, borrowed from the amsmath package, for use only before hyphens or dashes. It prevents the line break that is normally allowed after the following sequence of dashes.

Fixing font sizes

The new fix-cm package, by Walter Schmidt, changes the CM font definition (.fd) files so that similar design sizes are used in both the OT1 and T1 encodings.

Font encodings

A number of options have been added to the textcomp package, enabling only available glyphs to be used. Also, the ‘NFSS font families’ are now divided into five different groups according to the subset of glyphs each provides from the full collection of symbols in the TS1 encoding. Given sufficient information about a font family textcomp will use this in order to limit the typesetting to those glyphs that are available. Use of this mechanism has also enhanced \oldstylenums to use the current font if possible.

Displaying font tables

With the nfssfont package you can now specify the font to display by giving its ‘NFSS classification’, rather than needing to know its external font file’s name. It is also now possible to generate large collections of font tables in batch mode by providing a suitable input file.

New input encodings

The inputenc package has been extended as follows: macce input encoding (Apple Central European), thanks to Radek Tryc and Marcin Wolinski; cp1257 for Baltic languages; latin10, thanks to Ionel Ciobîcă. The euro symbol has by now been added to several encodings: ansinew, cp1250 and cp1252 (which also has another addition), whilst cp858 adds it to cp850.

Unicode input

Partial, experimental support for text files that use the Unicode encoding form UTF-8 is now provided by the option utf8 for the inputenc package.

The only Unicode text file characters supported by the current version are those based on the most common inputs for glyphs from the small collection of standard L^{A}TEX Latin encodings.

And finally … pict2e

The old, non-functional version of this package has been removed as there is now a fully working version from Hubert Gäßlein and Rolf Niepraschk. It is described in The L^{A}TEX Manual.

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 2003, all rights reserved.